# Digital requirements engineering with an INCOSE-derived SysML meta-model

### v1

@@ 3 Discussion of Benefits and Challenges
The immediate advantage of this MBSR approach was the ability to keep architecture modeling activities confined to a single tool, and without the loss of expressiveness that would normally result from using classical SysML Requirements. Report templates were created using the full MBSR expression, and according to stakeholder expectations for engineering reviews. The system architecture model and cohabitated requirements were exported to targeted stakeholder views, including custom PowerPoint templates for an architecture overview and a Requirement Set formal review, an internal website (Web Report), and Excel spreadsheets (adding relevant table columns as needed). While this targeted approach of INCOSE-derived MBSR usage received overwhelmingly positive feedback from the team, it came with its own challenges that became evident as limitations of the tool were encountered.
@@ 3.1 Observed Benefits
Observed benefits during the NASA/ESA/JPL joint Mars Sample Return project included: improved requirement statement quality due to the use of a common requirement Pattern, Rules, and accompanying Attributes; increased accuracy and assurance of model views using generated custom reports; access to system model elements for increased traceability and specificity; and model-assisted traceability views of Key Driving Requirements across multiple levels of the system hierarchy. Table 2 provides a more complete list with perceived impact on the project indicated.
@@ 3.2 Observed Challenges
Applying INCOSE-derived MBSR to a real project met with real challenges, primarily that the SysML tool in use is slow to generate reports, render views of model data and adorning legend items, and save model updates, causing significantly delayed cycle times during MBSR meta-modeling and modeling of the system. Other challenges of this approach relate to the SysML tool’s lack of RMT facilities that would automate some model updates and provide high assurance of their correctness. In this case, only one engineer was developing the requirements at the time, and formal change management had not been activated during this technology development phase. A more detailed list of the observed challenges and perceived impacts on the project are sorted in Table 3.
@@
Redefining |
@@
| Attributes with multiplicity ¿1 fails to copy/paste; paste chooses the first element with that name, not the same exact element copied | |
@@ 3.3 Applicability to Standard Guidance
This INCOSE-derived MBSR Profile supports the requirements engineering practices according to the NASA SE Handbook guidelines and Appendix C checklist (Hirshorn et al., 2017). Although the NASA SE Handbook does not present clearly defined characteristics, the glossary entry and Appendix C checklist contains language that may be mapped to all of the INCOSE GtWR Characteristics (refer to Table 1): (C1) ‘necessary …to meet mission and system goals and objectives’; (C2) ‘compliance’; (C3) ‘unambiguous in meaning’ and ‘complies with the project’s template and style rules’; (C4) ‘completeness’; (C5) ‘not redundant’; (C6) ‘feasible to obtain’; (C7) ‘verifiability/testability’; (C8) ‘correct’; (C9) ‘clear’ and ‘clarity’; (C10) ‘completeness’; (C11) ‘consistency’ and ‘not in conflict with one another’; (C12) ‘technically feasible’; (C13) ‘clarity’ and ‘adequately related with respect to terms used’; (C14) ‘can be validated’; and (C15) ‘correctness’ and ‘completeness’. By contrast, the ISO/IEC/IEEE (2018) standard provides definitions for 14 out of the 15 Characteristics and clearly categorizes them for individual requirements and for requirement sets. MBSR supports the bidirectional traceability and the creation of NASA-requested artifacts such as the Requirements Allocation Sheet, TBX report, and Requirements Verification and Validation Matrices. Further discussion of the relationships among the NASA SE Handbook, the ISO 29148:2018 standard, and the INCOSE GtWR may be found in the ontology section of the NRM. While the NASA SE Handbook and INCOSE GtWR (and related guides and manuals) are complementary, the GtWR provided enumerated and defined precision amenable to SysML meta-modeling and reuse in the system architecture model.
@@ 4 Related Work
Salado (2023) presents three approaches to model-based requirements found in the literature: 1) dedicated classes and flagged models, 2) math- and property-based models of requirements, and 3) semantic extensions to model the problem space. The relationship of MBSR to template-based textual requirements is discussed in Sec. 2.2 and standard SysML textual requirements (“dedicated classes”) are discussed in Sec. 1.3. This section provides a brief overview of related model-based approaches while noting similarities and differences with MBSR as presented in this paper.
@@ 4.1 System Models as Requirements
Wach and Salado (2022) presents a model-based method for capturing requirements using standard SysML elements and diagrams other than the textual-based SysML
AbstractRequirement
. As discussed in their paper, this method unnecessarily constrains the solution space and is therefore considered poor requirements engineering practice. Flagged models as requirements will be unfamiliar to stakeholders and may cause significant confusion compared to “shall” statements (Wheatcraft et al., 2022). MBSR extends the use of textual statements with deeper connectivity to the system model compared to standard SysML, without relying on SysML diagrams to model the problem space. Statement patterns are used in MBSR both to assist readability and to facilitate model definition and traceability. While SysML elements are used in MBSR slots, they may remain placeholders until design activities commence and more details of the solution space are known.
@@ 4.2 Mathematical Models of Requirements
Wymorian theory of requirements is based in set theory (Wymore, 1993) and is not directly applicable to SysML models. This formulation of the problem space enables mathematical queries to assist in requirements V&V and may in particular assess completeness of a requirement set. The basic modeling construct is called a system design requirement and is defined as a sextuple including: 1) input/output requirement, 2) technology requirement, 3) performance requirement, 4) cost requirement, 5) trade-off requirement, and 6) system test requirement. Compared to Wymorian theory of requirements, MBSR is more relevant to modern MBSE practice because of its integration with SysML models. Integration of this alternative MBSE theory with SysML remains an open area of research (Wach and Salado, 2019), and SysML v2 may provide new research opportunities due to its strong semantic foundations compared to v1.
@@
Property-Based Requirements (PBRs) are another mathematical formulation of system requirements based on a semilattice (Micouin, 2008; Bernard, 2012). This formulation may be applied to other modeling languages such as AADL, Modelica, and VHDL-AMS, but Micouin (2008) focuses the presentation on SysML. Like MBSR, PBR extends the standard SysML requirement element by creating a PBRequirement stereotype, and like MBSR, the PBRequirement stereotype provides four primary attributes representing 1) Condition, 2) Carrier, 3) Property, and 4) Domain (Micouin, 2008). What is considered a well-formed PBR refers to its completion of these attributes to form a mathematical constraint, which in light of Wheatcraft and Ryan (2023) is insufficient for mature requirements engineering practice. PBR does not distinguish between stakeholder needs and system requirements, whereas INCOSE-derived MBSR support for needs and need sets is in development. Composite requirements, hierarchical nesting of requirements, and automated identification of dependencies are features of PBR that MBSR supports using SysML (e.g. Fig. 7). OMG (2022) presents PBR as a non-normative extension with examples in Annex E, but this formulation differs from Micouin (2008) in that it focuses on “quantitative specification of numerical parameters, relationships, equations and/or constraints.” PBRs, as presented in the SysML v1.7 specification Annex E.8, may be adapted to the MBSR Profile to provide formal verification capabilities especially suited for numerical- and logic-focused requirements while benefitting from the INCOSE GtWR ontology and rules. Lu et al. (2008) presents another property-based formulation of requirements using a custom object-oriented modeling tool, and discusses similar benefits to MBSR, such as enhanced traceability, consistency, completeness, maintainability, and integration with artifact and report generation.
@@
Ontology-based requirements engineering is an active area of MBSE research with promising results that may contribute to the ASoT with reusable ontology-defined terms (Avdeenko and Pustovalova, 2016; Yang et al., 2019; Lorch et al., 2024). An ontology-based requirement may be composed of slots referring to Pattern Slots of the Requirement Statement, and the Attributes contributing to the complete Requirement Expression. Through establishing ontology relationships and axioms, model-based requirements V&V may be assisted with tool automation ensuring (C10) completeness, (C11) consistency, (C15) correctness, (C13) “unambiguity”, and (C14) “traceability” (Avdeenko and Pustovalova, 2016). While the INCOSE-derived MBSR Profile may appear to implement an ontology, it lacks the formal semantics defined by ontology languages such as the Web Ontology Language (OWL) (World Wide Web Consortium, 2012). Therefore, the MBSR Profile is not currently capable of being exported to a standard ontology format, although a mapping from the meta-model to an ontology may be technically feasible. Integration with ontology tools may enhance the reasoning capability of a SysML model, but at the cost of tooling complexity, which MBSR attempts to avoid.
@@ 4.3 Semantic Extensions to Model the Problem Space
Alternatively, semantic extensions to SysML may be employed to model the problem space, known as True Model-Based Requirements (TMBR) (Salado and Wach, 2019). TMBR derives from the Wymorian theory of requirements (Sec. 4.2) and so “attempts to model any type of requirement as a set of (or sets) of required input/output transformations” (Salado, 2023). TMBR is not intended for modeling stakeholder needs (Salado, 2023), compared to INCOSE-derived MBSR, which includes Needs and Need Sets in its meta-model. The set theory foundations of TMBR prevent “formal flaws in problem formulation, such as enforcing design solutions or leaving the requirement unbounded” (Salado, 2023), corresponding to Characteristics (C15) Correct, (C12) Feasible and (C14) Able to be validated. Due to the reuse of SysML features to model the problem space, stakeholders may find TMBR difficult to understand compared to traditional requirement statements (Salado and Shadab, 2024). TMBR and MBSR are similar in that they use SysML as the primary modeling language and make use of SysML attributes with defined quantities and constraints.
5 Limitations and Future Work
@@
Limitations of the present work are primarily due to scoping and available SysML v1 tool capabilities. Although experimental design as done by Salado and Shadab (2024) to compare this MBSR approach with other tool-enabled requirements engineering methods was not conducted, qualitative validation of this MBSR approach was conducted in the context of two real-world system development projects at NASA Jet Propulsion Laboratory by the first author. Producing experimental evidence of the MBSR contribution toward satisfying the 15 Characteristics of Well-formed Requirements and Sets of Requirements (Table 1) is left as future work; readers are encouraged to refer to Wheatcraft and Ryan (2023) and accompanying manuals available for free to INCOSE members. Contractual agreements currently limit the exposure of proprietary information such as the roughly 500 MBSRs written, so a minimal example of an asteroid sample collection spacecraft was used for demonstration. Although included in the MBSR Profile and linked to the applicable INCOSE GtWR Attribute sets, Needs and Need Sets were not used and evaluated in the projects, and their model-based patterns are currently undefined. We provide the MBSR SysML Profile in an open-source online repository (The Authors, 2024) so other organizations and researchers may adapt it and run experiments to further the systems engineering community’s understanding of the effectiveness of this approach.
Future work will continue to explore the feasibility and effectiveness of INCOSE-derived MBSR.
Some INCOSE GtWR Rules, such as those disallowing certain words and phrases, may be automatically checked using simple string matching, achieving similar capability to existing requirements quality management tools.
Such a capability may be encoded in automated validation suites using custom scripts similar to the ones we wrote for metrics and legends, although the potential performance impact is a concern addressable by future research.
Macros may be created that employ automated validation suite results to add and remove
Satisfy
and
Violate
relationships for Rules amenable to automation.
@@
Automatically-generated requirement statements, as proposed by Herber et al. (2022), are provided as a read-only derived attribute, but future work is needed to make this ‘Derived Text’ attribute more readable while not interfering with model element naming conventions: such as by adding a stereotype to MBSR-linked elements that provides text attributes for defining the requirement statement fragment depending on its intended pattern slot.
The Condition pattern slot may be further decomposed following Mavin and Wilkinson (2010), which categorizes conditions into event-driven, unwanted behavior, state-driven, and optional feature; or it may be decomposed according to a hierarchical ontology as shown in Carson (2015) and the GtWR Appendix C.
The capacity of an ASoT with MBSRs to answer targeted mission programmatic questions is an area of future research that depends on the high maturity of both the system architecture model and requirements engineering organizational processes.
@@
The revised semantics in SysML v2 present an opportunity to adapt the MBSR Profile to the new language and to research the change impact and potential benefits. SysML v2 uniformly applies the modeling language design pattern of definition and usage to requirements; it defines requirements necessarily as constraints with boolean satisfiability; stakeholder concerns are more explicitly modeled compared to SysML v1; and metadata elements are used instead of stereotypes. SysML v2 also provides a standardized API for accessing the system architecture model which may ease future research in integrating MBSR with NLP-based and formal requirements engineering tools. This shift toward more model-based requirements presents new opportunities to enrich the ASoT with metadata and automated model validation derived from the INCOSE GtWR.
6 Conclusions

### v2

@@ 4 Discussion
The immediate advantage of this MBSR approach was the ability to keep architecture modeling activities confined to a single tool, and without the loss of expressiveness that would normally result from using classical SysML Requirements. Report templates were created using the full MBSR expression, and according to stakeholder expectations for engineering reviews. The system architecture model and cohabitated requirements were exported to targeted stakeholder views, including custom PowerPoint templates for an architecture overview and a Requirement Set formal review, an internal website (Web Report), and Excel spreadsheets (adding relevant table columns as needed). While this targeted approach of INCOSE-derived MBSR usage received overwhelmingly positive feedback from the team, it came with its own challenges that became evident as limitations of the tool were encountered (Sec. 4.2).
@@ 4.1 Observed Benefits
Observed benefits compared to unstructured document-based and classical SysML requirements during the NASA / ESA / JPL joint Mars Sample Return project included: improved requirement statement quality due to the use of a common requirement Pattern, Rules, and accompanying Attributes; increased accuracy and assurance of model views using generated custom reports; access to system model elements for increased traceability and specificity; and model-assisted traceability views of Key Driving Requirements across multiple levels of the system hierarchy. Table 1 provides a more complete list with perceived impact on the project during Pre-Phase A.
@@ 4.2 Observed Challenges
Applying INCOSE-derived MBSR to a real project met with real challenges (Table 2), primarily that the SysML tool in use was slow to generate reports, render views of model data and adorning legend items, and save model updates, causing significantly delayed cycle times during MBSR meta-modeling and modeling of the system. Other challenges of this approach relate to the SysML tool’s lack of RMT facilities that would automate some model updates and provide high assurance of their correctness. In this case, only one engineer was developing the requirements at the time, and formal change management had not been activated during this technology development phase. A more detailed list of the observed challenges and perceived impacts on the project are sorted in Table 2.
@@
| Redefining «AbstractRequirement» Text attribute breaks automatic underlining of quantity relations | |
@@
| Attributes with multiplicity >1 fails to copy/paste; paste chooses the first element with that name, not the same exact element copied | |
| Element naming conflicts interfere with the use of copy/paste in tables | |
@@ 4.3 Applicability to Standard Guidance
This INCOSE-derived MBSR Profile supports the requirements engineering practices according to the NASA SE Handbook guidelines and Appendix C checklist (NASA_SE_HB2017). Although the NASA SE Handbook does not present clearly defined characteristics, the glossary entry and Appendix C checklist contains language that may be mapped to all of the INCOSE GtWR Characteristics (refer to Table 3): (C1) ‘necessary …to meet mission and system goals and objectives’; (C2) ‘compliance’; (C3) ‘unambiguous in meaning’ and ‘complies with the project’s template and style rules’; (C4) ‘completeness’; (C5) ‘not redundant’; (C6) ‘feasible to obtain’; (C7) ‘verifiability/testability’; (C8) ‘correct’; (C9) ‘clear’ and ‘clarity’; (C10) ‘completeness’; (C11) ‘consistency’ and ‘not in conflict with one another’; (C12) ‘technically feasible’; (C13) ‘clarity’ and ‘adequately related with respect to terms used’; (C14) ‘can be validated’; and (C15) ‘correctness’ and ‘completeness’. By contrast, the ISO29148 standard provides definitions for 14 out of the 15 Characteristics and clearly categorizes them for individual requirements and for requirement sets. MBSR supports the bidirectional traceability and the creation of NASA-requested artifacts such as the Requirements Allocation Sheet, TBX report, and Requirements Verification and Validation Matrices. Further discussion of the relationships among the NASA SE Handbook, the ISO 29148:2018 standard, and the INCOSE GtWR may be found in the ontology section of the NRM. While the NASA SE Handbook and INCOSE GtWR (and related guides and manuals) are complementary, the GtWR provided enumerated and defined precision amenable to SysML meta-modeling and reuse in the system architecture model.
| ID | Name | Applicability | Derivation | NASA? | ISO? |
| C1 | Necessary | Needs & Requirements | Formal Transformation | ||
| C2 | Appropriate | Needs & Requirements | Formal Transformation | ||
| C3 | Unambiguous | Needs & Requirements | Agreed-to Obligation | ||
| C4 | Complete | Needs & Requirements | Agreed-to Obligation | ||
| C5 | Singular | Needs & Requirements | Formal Transformation | ||
| C6 | Feasible | Needs & Requirements | Agreed-to Obligation | ||
| C7 | Verifiable | Needs & Requirements | Agreed-to Obligation | ||
| C8 | Correct | Needs & Requirements | Formal Transformation | ||
| C9 | Conforming | Needs & Requirements | Formal Transformation | ||
| C15 | Correct | Need Sets & Requirement Sets | Formal Transformation |
@@ 5 Related Work
salado2023model presents three approaches to model-based requirements found in the literature: 1) dedicated classes and flagged models, 2) math- and property-based models of requirements, and 3) semantic extensions to model the problem space. The relationship of MBSR to template-based textual requirements is discussed in Sec. 2.2, and classical SysML requirements (“dedicated classes”) are discussed in Sec. 1.3. This section provides a brief overview of related model-based requirement approaches while noting similarities and differences with MBSR as presented in this paper.
@@ 5.1 System Models as Requirements
wach2022need presents a model-based method for capturing requirements using standard SysML elements and diagrams other than the textual-based SysML «AbstractRequirement». As discussed in their paper, this method unnecessarily constrains the solution space and is therefore considered poor requirements engineering practice. Flagged models as requirements will be unfamiliar to stakeholders and may cause significant confusion compared to “shall” statements (NRM2022). MBSR extends the use of textual statements with deeper connectivity to the system model compared to standard SysML, without relying on SysML diagrams to model the problem space. Statement patterns are used in MBSR both to assist readability and to facilitate model definition and traceability. While SysML elements are used in MBSR slots, they may remain placeholders until design activities commence and more details of the solution space are known.
@@ 5.2 Mathematical Models of Requirements
Wymorian theory of requirements is based in set theory (wymore1993model) and is not directly applicable to SysML models. This formulation of the problem space enables mathematical queries to assist in requirements V&V and may in particular assess completeness of a requirement set. The basic modeling construct is called a system design requirement and is defined as a sextuple including: 1) input/output requirement, 2) technology requirement, 3) performance requirement, 4) cost requirement, 5) trade-off requirement, and 6) system test requirement. Compared to Wymorian theory of requirements, MBSR is more relevant to modern MBSE practice because of its integration with SysML models. Integration of this alternative MBSE theory with SysML remains an open area of research (wach2019can), and SysML v2 may provide new research opportunities due to its strong semantic foundations compared to v1.
@@
Property-Based Requirements (PBRs) are another mathematical formulation of system requirements based on a semilattice (micouin2008toward; bernard2012requirements). This formulation may be applied to other modeling languages such as AADL, Modelica, and VHDL-AMS, but micouin2008toward focuses the presentation on SysML. While PBRs focus on the integration and computation of physical properties to prepare for system V&V, our work emphasizes the importance of well-formed requirement statements whose components are SysML model elements. Like MBSR, PBR extends the standard SysML requirement element by creating a PBRequirement stereotype, and like MBSR, the PBRequirement stereotype provides four primary attributes representing 1) Condition, 2) Carrier, 3) Property, and 4) Domain (micouin2008toward). What is considered a well-formed PBR refers to its completion of these attributes to form a mathematical constraint, which in light of GtWR2023 is insufficient for mature requirements engineering practice. PBR does not distinguish between stakeholder needs and system requirements, whereas INCOSE-derived MBSR supports needs and need sets in addition to requirement expressions and requirement sets. Composite requirements, hierarchical nesting of requirements, and automated identification of dependencies are features of PBR that MBSR supports using SysML (e.g. Fig. 3). SysML1.7 presents PBR as a non-normative extension with examples in Annex E, but this formulation differs from micouin2008toward in that it focuses on “quantitative specification of numerical parameters, relationships, equations and/or constraints.” PBRs, as presented in the SysML v1.7 specification Annex E.8, may be adapted to the MBSR Profile to provide formal verification capabilities especially suited for numerical- and logic-focused requirements while benefitting from the INCOSE GtWR ontology and rules. lu2008requirement presents another property-based formulation of requirements using a custom object-oriented modeling tool, and discusses similar benefits to MBSR, such as enhanced traceability, consistency, completeness, maintainability, and integration with artifact and report generation.
@@
Ontology-based requirements engineering is an active area of MBSE research with promising results that may contribute to the ASoT with reusable ontology-defined terms (avdeenko2016ontology; yang2019ontology; lorch2024formal). An ontology-based requirement may be composed of slots referring to Pattern Slots of the Requirement Statement, and the Attributes contributing to the complete Requirement Expression. Through establishing ontology relationships and axioms, model-based requirements V&V may be assisted with tool automation ensuring (C10) completeness, (C11) consistency, (C15) correctness, (C13) “unambiguity”, and (C14) “traceability” (avdeenko2016ontology). While the INCOSE-derived MBSR Profile may appear to implement an ontology, it lacks the formal semantics defined by ontology languages such as the Web Ontology Language (OWL) (OWL2). Therefore, the MBSR Profile is not currently capable of being exported to a standard ontology format, although a mapping from the meta-model to an ontology may be technically feasible. Integration with ontology tools may enhance the reasoning capability of a SysML model, but at the cost of tooling complexity, which MBSR attempts to avoid.
@@ 5.3 Semantic Extensions to Model the Problem Space
Alternatively, semantic extensions to SysML may be employed to model the problem space, known as True Model-Based Requirements (TMBR) (salado2019constructing). TMBR derives from the Wymorian theory of requirements (Sec. 5.2) and so “attempts to model any type of requirement as a set of (or sets) of required input/output transformations” (salado2023model). TMBR is not intended for modeling stakeholder needs (salado2023model), compared to INCOSE-derived MBSR which includes Needs and Need Sets in its meta-model (Fig. 4). The set theory foundations of TMBR prevent “formal flaws in problem formulation, such as enforcing design solutions or leaving the requirement unbounded” (salado2023model), corresponding to Characteristics (C15) Correct, (C12) Feasible and (C14) Able to be validated. Due to the reuse of SysML features to model the problem space, stakeholders may find TMBR difficult to understand compared to traditional requirement statements (salado2024comparative). TMBR and MBSR are similar in that they use SysML as the primary modeling language and make use of SysML attributes with defined quantities and constraints. As a language extension, MBSR is compatible with other SysML profiles including SYSMOD (weilkiens2020sysmod).
6 Conclusion
While it is too early to determine the economic value of using INCOSE-derived MBSRs compared to traditional requirements engineering methods, initial results indicate the technique had a positive impact on a real project. We now present limitations of the present work, highlight paths forward for future work, and conclude with summarizing remarks.
6.1 Limitations
@@
Limitations of the present work are primarily due to scoping and available SysML v1 tool capabilities. While our experiments have primarily made use of Cameo Systems Modeler, we posit that non-normative extensions to SysML v1 such as tables, matrices, and scripting capabilities are not restricted to certain tools, and that other tools will work in similar ways that we cannot fully describe here. Therefore, the MBSR technique is presented as a SysML extension, rather than a tool-grafted technique. Although experimental design as done by salado2024comparative to compare this MBSR approach with other tool-enabled requirements engineering methods was not conducted, qualitative validation of this MBSR approach was conducted in the context of two real-world system development projects at NASA Jet Propulsion Laboratory by the first author. Producing experimental evidence of the MBSR contribution toward satisfying the 15 Characteristics of Well-formed Requirements and Sets of Requirements (Table 3) is left as future work; readers are encouraged to refer to GtWR2023 and accompanying manuals available for free to INCOSE members. Contractual agreements currently limit the exposure of proprietary information such as the roughly 500 MBSRs written, so a minimal example of an asteroid sample collection spacecraft was used for demonstration. Although included in the MBSR Profile and linked to the applicable INCOSE GtWR Attribute sets, Needs and Need Sets were not used and evaluated in the projects. We provide the MBSR SysML profile in an open-source online repository (herber2024repo) so other organizations and researchers may adapt it and run experiments to further the systems engineering community’s understanding of the effectiveness of this approach.
6.2 Future Work
Future work will continue to explore the feasibility and effectiveness of INCOSE-derived MBSR. Future work remains to make full use of SysML modeling tool capabilities in conjunction with other tools such as MATLAB, Modelica, and Python to exploit the SysML-based traceability that the MBSR technique provides. Some INCOSE GtWR Rules, such as those disallowing certain words and phrases, may be automatically checked using simple string matching, achieving similar capability to existing requirements quality management tools. Such a capability may be encoded in automated validation suites using custom scripts similar to the ones we wrote for metrics and legends, although the potential performance impact is a concern addressable by future research. Macros may be created that employ automated validation suite results to add and remove «Satisfy» and «Violate» relationships for Rules amenable to automation. Many other model validation rules may be added that further enhance the model-based requirements V&V automation afforded by MBSR, such as checking that architecture decomposition levels match across MBSR-referenced model elements, or checking for conflicting or redundant requirements.
@@ 6.2 Future Work
Automatically-generated requirement statements, as proposed by herber2022model, are provided as a read-only derived attribute, but future work is needed to make this ‘Derived Requirement Statement’ attribute more readable while not interfering with model element naming conventions: such as by adding a stereotype to MBSR-linked elements that provides text attributes for defining the requirement statement fragment depending on its intended pattern slot. Names of the Pattern Slots and their element values could be used with Large Language Models or other advanced Natural Language Processing models to generate natural language sentences with proper syntax and grammar. The Condition pattern slot may be further decomposed following mavin2010big, which categorizes conditions into event-driven, unwanted behavior, state-driven, and optional feature; or it may be decomposed according to a hierarchical ontology as shown in carson2015implementing and the GtWR Appendix C. The capacity of an ASoT with MBSRs to answer targeted mission programmatic questions is an area of future research that depends on the high maturity of both the system architecture model and requirements engineering organizational processes.
@@
The revised semantics in SysML v2 present an opportunity to adapt the MBSR Profile to the new language and to research the change impact and potential benefits. SysML v2 uniformly applies the modeling language design pattern of definition and usage to requirements; it defines requirements necessarily as constraints with boolean satisfiability; stakeholder concerns are more explicitly modeled compared to SysML v1; and metadata elements are used instead of stereotypes. Furthermore, SysML v2 specifies a textual syntax in addition to the graphical syntax, potentially making model representations such as Fig. 3.3 more readable and amenable to automation. SysML v2 also provides a standardized API for accessing the system architecture model, which may ease future research in integrating MBSR with NLP-based and formal requirements engineering tools. This shift toward more model-based requirements presents new opportunities to enrich the ASoT with metadata and automated model validation derived from the INCOSE GtWR.
6.3 Summary